The ratio of the width to the length of a rectangle is 5:6. The perimeter of the rectangle is 121 inches. Find the area of the rectangle
Answer by stanbon(75887)

The ratio of the width to the length of a rectangle is 5:6. The perimeter of the rectangle is 121 inches. Find the area of the rectangle
----
width = 5x
length = 6x
-----
Perimeter = 2(length + width)
121 = 2(6x+5x)
121 = 2(11x)
11 = 2x
x = 5.5
------
width = 5(5.5) inches
length = 6(5.5) inches
-----
Area = length * width
Area = 30*5.5^2 = 907.5 sq.in.
-------------------
Cheers,
Stan H.
-------------------